OpenStudy (anonymous):

in triangle DEF, sides DE and DF are both 6 feet 3 inches long. This length is 0.75 times the length of side FE. What type of triangle is DEF and what is its perimeter?

OpenStudy (anonymous):

DE=DF=75 inches 75=0.75*FE=>FE=100 INCHES The triangle is isosceles since it has 2 equal sides P=100+75+75=250 inches=20 feet 10 inches i hope i did the conversions right,am not that good with inches